### v1

Abstract.
In this work, we develop the method of multipliers for electromagnetic Dirac operators and establish sufficient conditions on the magnetic and electric fields that guarantee the absence of point spectrum. Our results can be viewed as a generalization of those obtained by Cossetti, Fanelli and Krejčiřík (2020), which treated the purely magnetic case. We also emphasize that, in the massless case, our approach covers Coulomb-type potentials of the form $V(x)=\frac{1}{|x|}\big{(}\nu\mathbb{I}+\mu\beta+i\delta\beta\big{(}\boldsymbol{\alpha}\cdot\frac{x}{|x|}\big{)}\big{)}.$
1. Introduction
In this work we are interested in spectral properties of the electromagnetic Dirac operator
$H_{m}(A,V)=-i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}+m\beta+V.$ (1.1)
Here $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3}),$ with $\alpha_{j},$ $j=1,2,3$ and $\beta$ being the standard $4\times 4$ Hermitian Dirac matrices, $m\geq 0$ denotes the mass of the particle, $V$ is a potential function that associates to each $x\in\mathbb{R}^{3}$ a $4\times 4$ Hermitian matrix $V(x).$ Moreover $\nabla_{\!\text{A}}$ stands for the magnetic gradient $\nabla-iA,$ where $A\colon\mathbb{R}^{3}\to\mathbb{R}^{3}$ is a magnetic vector potential. The operator (1.1) represents the Hamiltonian of a relativistic particle of spin-$\tfrac{1}{2}$ under the action of an electric field $V$ and a magnetic field $B=\nabla\times A.$
Compared to the cases of purely electric [23, 28, 25, 10, 15, 32, 33] or purely magnetic [2, 19, 18, 6, 22] perturbations, the literature available related to the absence of point spectrum of electromagnetic Dirac operators is much scarcer and much less complete. The results presented in this work are motivated by the need to fill in this gap. One of the first works available on this topic is due to Roze [31], who considered electric potentials of the form $V(x)=v(x)\mathbb{I},$ and proved that if $A$ and $v$ are smooth and satisfy
$\lim_{|x|\to\infty}|x|(|A(x)|+|v(x)|)=0,$ (1.2)
then the Dirac operator has no eigenvalues embedded in the continuous spectrum. A similar result was proved by Kalf in [24] where he replaced the decay assumption (1.2) by an analogous assumption involving the magnetic field $B$ instead of the magnetic potential $A$. Thus giving a gauge invariant condition for the absence of eigenvalues. The result of Kalf was later extended to more general matrix-valued potentials in [3]. In a more recent result [30] the authors investigated on spectral properties of Dirac operators with magnetic fields of constant direction and Coulomb-type electric potentials.
In the present paper, we will develop the so-called method of multipliers in a general setting in order to establish a source of sufficient conditions (decay/Hardy-type) on the potentials to guarantee the absence of point spectrum (including possibly embedded eigenvalues) of the Dirac Hamiltonian $H_{m}(A,V)$ in dimension $d\geq 3.$ See [8] for a recent survey on the method of multipliers in spectral theory.
In the last decade, starting from the pioneering work [17], this method was shown to be a very powerful tool in spectral theory and, in particular, in proving absence of eigenvalues (embedded and discrete) of very different Hamiltonians [6, 9, 8, 7, 16, 17, 1, 22]. Despite its robustness, this method does not seem to apply directly to general Dirac operators. Indeed, some technical difficulties arise due to the lack of positivity of certain commutators, which is also related to the lack of lower semi-boundedness of the operator itself. See references  [11, 12, 14, 15, 21, 27] for related results obtained using alternative techniques. To recover such positivity, a natural strategy is to employ the well-known supersymmetric structure of the free Dirac operator, $H_{m}(0,0)=-i\boldsymbol{\alpha}\cdot\nabla+m\beta$. This means that squaring out the operator gives
$H_{m}(0,0)^{2}=\begin{pmatrix}H_{P}+m^{2}\mathbb{I}_{\mathbb{C}^{2}}&0\\ 0&H_{P}+m^{2}\mathbb{I}_{\mathbb{C}^{2}},\end{pmatrix}$
where $H_{P}:=-\Delta\mathbb{I}_{\mathbb{C}^{2}}+\boldsymbol{\sigma}\cdot B$ denotes the Pauli operator in $L^{2}(\mathbb{R}^{3})^{2}.$ Here $\sigma=(\sigma_{1},\sigma_{2},\sigma_{3})$ are the standard $2\times 2$ Hermitian Pauli matrices. An important consequence of this spectral relation is that proving absence of eigenvalues of $H_{m}(0,0)$ reduces to prove absence of eigenvalues of second order operators of Schrödinger type for which the aforementioned positivity is guaranteed. It is easy to see that the supersymmetric property is maintained for a purely magnetic Dirac operator, namely for $H_{m}(A,0)$ (see [8] and [22]). However, introducing a general electric potential $V$ breaks this property, making the analysis of such operators much harder. Overcoming this difficulty is precisely the aim of the present work.
The first result of this paper shows that, if the potential satisfies a suitable Hardy-type smallness condition, no eigenvalues are created in the spectral gap.
Theorem 1.1.
Let $d\geq 3$ and let $V:\mathbb{R}^{d}\to\mathbb{C}^{N\times N},$ $N:=2^{\lfloor(d+1)/2\rfloor},$ be such that the following inequality
$\|V\psi\|\leq\|H_{m}(A)\psi\|-m\|\psi\|$ (1.3)
holds for every $\psi\in H^{1}(\mathbb{R}^{d})$. Then $\sigma_{p}(H_{m}(A,V))\cap(-m,m)=\varnothing$.
Remark 1.1.
It follows from an easy computation that
$\|H_{m}(A)\psi\|-m\|\psi\|\leq\|-i(\alpha\cdot\nabla_{\!\text{A}})\psi\|.$
Thus, condition (1.3) can be interpreted as a requirement for the potential $V$ to satisfy a magnetic Hardy inequality. As with Schrödinger-type operators, this kind of condition is not usually sufficient to guarantee the total absence of eigenvalues; it only ensures the absence of discrete eigenvalues.
The main result of this paper reads as follows.
Theorem 1.2.
Let $d\geq 3.$ Suppose that $V:\mathbb{R}^{d}\to\mathbb{C}^{N\times N}$ is such that $V\in C^{\infty}(\mathbb{R}^{d}\setminus\{0\})$ and that $V$ is bounded outside any ball containing the origin. Additionally, assume that there exist constants $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4}\geq 0$ satisfying
$\left(\frac{4d-6}{d-2}\right)\varepsilon_{1}+\left(\frac{4}{d-2}\right)\varepsilon_{2}\,\varepsilon_{3}+2m\varepsilon_{4}^{2}+\left(\frac{8d-8}{d-2}\right)\varepsilon_{3}+\left(\frac{4d-4}{d-2}\right)\varepsilon_{2}\leq 2,$ (1.4)
and assume that
$\int|x|^{2}|B|^{2}|\psi|^{2}$ $\leq\varepsilon_{1}^{2}\|\nabla_{\!\text{A}}\psi\|^{2},$ (1.5)
$\int|x|^{2}|\nabla V|^{2}|\psi|^{2}$ $\leq\varepsilon_{2}^{2}\|\nabla_{\!\text{A}}\psi\|^{2},$ (1.6)
$\sup_{x\in\mathbb{R}^{d}}|x||V(x)|$ $\leq\varepsilon_{3},$ (1.7)
and, if $m\neq 0$ or $\{\beta,V\}\neq 0,$ assume further that
$\int|x||\nabla V||\psi|^{2}\leq\varepsilon_{4}^{2}\|\nabla_{\!\text{A}}\psi\|^{2}$ (1.8)
hold true for every $\psi\in H^{1}(\mathbb{R}^{d})$. Then $\sigma_{p}(H_{m}(A,V))=\varnothing$.
Theorem 1.2 remains valid if we move the singularity of the potentials from the origin to another point in $\mathbb{R}^{d}$.
Corollary 1.1.
Let $d\geq 3.$ Suppose that $V:\mathbb{R}^{d}\to\mathbb{C}^{N\times N}$ is such that $V\in C^{\infty}(\mathbb{R}^{d}\setminus\{x_{0}\})$ and that $V$ is bounded outside any ball containing the origin. If $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4}\geq 0$ satisfy (1.4) and $B$ and $V$ satisfy (1.5)-(1.8), then $\sigma_{p}(H_{m}(A,V))=\varnothing.$
Another easy consequence of Theorem 1.2 is that as soon as this theorem is valid for an electric potential $V$ it is also valid for any conjugation of $V$ by unitary matrices.
Corollary 1.2.
Let $d\geq 3.$ Suppose that $V$ is as in Theorem 1.2. If $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4}$ satisfy (1.4) and $B$ and $V$ satisfy (1.5)-(1.8), then for any unitary matrix $P\in\mathbb{C}^{N\times N}$ one has $\sigma_{p}(H_{m}(A,PVP^{\ast}))=\varnothing,$
Remark 1.2 (Decay behavior of $B$ and $V$).
We observe that the Hardy-type conditions (1.5), (1.6) and (1.8) prescribe a behavior at infinity of the magnetic field $B$ and the electric field $V:$ By virtue of the standard Hardy-inequality
$\int\frac{|\psi|^{2}}{|x|^{2}}\leq\frac{4}{(d-2)^{2}}\int|\nabla\psi|^{2},\qquad\forall\,\psi\in H^{1}(\mathbb{R}^{d}),$ (1.9)
if $W$ satisfies the pointwise condition $|W|\leq\mu|x|^{-2}$ for some $\mu>0,$ then
$\int|W||\psi|^{2}\leq\frac{4\mu}{(d-2)^{2}}\int|\nabla\psi|^{2}=:c_{\mu,d}\int|\nabla\psi|^{2}.$ (1.10)
This provides a sufficient condition for assumptions (1.5), (1.6) and (1.8) to be satisfied, i.e. $|W|\approx\mu|x|^{-2},$ for some suitably small coupling constant $\mu>0.$ Notice that in the specific cases, namely (1.5), (1.6) and (1.8), one has $|W|=|x|^{2}|B|^{2},$ $|W|=|x|^{2}|\nabla V|^{2},$ and $|W|=|x||\nabla V|,$ respectively.
On the other hand, one can prove (see [20, Thm. 2.1] and [35]) that if $W$ satisfies a Hardy-type inequality (1.10), then there exist $c,R>0$ such that
$|W(x)|\leq\frac{c}{|x|^{2}},\quad|x|>R.$
So the $|x|^{-2}$ decay behavior is also a necessary condition for the validity of (1.5), (1.6) and (1.8).
Finally observe that, as a consequence of the diamagnetic inequality (see, for instance, [26])
$\left|\nabla|\psi(x)|\right|\leq|\nabla_{\!\text{A}}\psi(x)|,$
it is enough to check the validity of (1.5), (1.6) and (1.8) above with the magnetic gradient replaced by the standard gradient.
Remark 1.3 (Refined version of condition (1.8)).
As one can see tracing the proof of Theorem 1.2, condition (1.8) can be relaxed in the following way depending on whether we are interested in the behavior close to the origin or at infinity:
$\int\langle x\rangle|\nabla V||\psi|^{2}\leq\varepsilon_{4}^{2}\|\nabla_{\!\text{A}}\psi\|^{2},$ (1.11)
here $\langle x\rangle:=(1+|x|^{2})^{1/2}$ is the standard notation for Japanese brackets. The alternative condition (1.11) shows that in the massive case, or if ${\beta,V}\neq 0$, we can indeed allow for Coulomb-type decay near the origin, but not at infinity.
Remark 1.4 (Coulomb-type potentials).
Notice that in the massless case, i.e. $m=0,$ since (1.8) is no longer needed, Theorem 1.2 also covers Coulomb-type potentials, that is potentials of the form
$V(x)=\frac{\nu}{|x|}\mathbb{V},$ (1.12)
for sufficiently small $\nu$ (regardless of its sign) and where $\mathbb{V}$ is any fixed, constant-coefficients, Hermitian matrix. In particular, the following special case
$\mathbb{V}=\mathbb{V}_{\text{el}}+\mathbb{V}_{\text{sc}}+\mathbb{V}_{\text{am}}:=\nu\mathbb{I}+\mu\beta+i\delta\beta\Big{(}\boldsymbol{\alpha}\cdot\frac{x}{|x|}\Big{)}$ (1.13)
is included, where the potentials $\mathbb{V}_{\text{el}},\mathbb{V}_{\text{sc}},\mathbb{V}_{\text{am}}$ are customarily called electric, scalar and anomalous magnetic, respectively and which are known to be perturbations that leave invariant the partial wave subspaces.
We emphasize that the impossibility to cover Coulomb-type potentials (1.12) in the massive case is not just technical, indeed it is known that, as soon as $m\neq 0,$ there are Coulomb-type perturbations that do produce discrete eigenvalues in the spectral gap (m,m)(-m,m) (see [13] and references therein for the case $A=0$ and $\mathbb{V}_{\text{el}}$ and [34] for the case $\mathbb{V}_{\text{sc}}$). The situation is different for anomalous magnetic potentials $\mathbb{V}_{\text{am}}.$ Indeed, in this case, since $\{\beta,\mathbb{V}_{\text{am}}\}=0,$ assumption (1.8) is no longer needed, thus Theorem 1.2 is valid for $V$ as in (1.12) with $\mathbb{V}=\mathbb{V}_{\text{am}}$ also in the massive case, namely for any $m\geq 0.$ We emphasize that this does not contradict the explicit eigenvalues and eigenfunctions obtained in [4] for Dirac operators with purely anomalous magnetic perturbations. Indeed, as can be seen from the proof of their result, if the coupling constant $\delta$ (there called $\lambda$) is sufficiently small, then the corresponding eigenfunctions found there are not $L^{2}$-integrable.
We are now interested in stating the corresponding results to Theorem 1.2 for the following potential
$V=V_{\text{el}}+V_{\text{sc}}+V_{\text{am}}:=v_{\text{el}}(x)\mathbb{I}+v_{\text{sc}}(x)\beta+i\beta(\boldsymbol{\alpha}\cdot\nabla)\phi_{\text{am}}(x).$
This is a generalization of the special case (1.12) with (1.13), and has been extensively studied as it is an interesting model in quantum mechanics (see, for example [4] for further information).
As we pointed out in Remark 1.4, Theorem 1.2 does apply to these potentials if $V_{\text{el}},V_{\text{sc}}$ and $V_{\text{am}}$ (or $v_{\text{el}},v_{\text{sc}}$ and $\phi_{\text{am}}$) satisfy its hypotheses. However, in these specific situations, stronger results can be obtained compared to the general case. This is why we will state and prove the corresponding results separately.
Theorem 1.3 (Massive Dirac operator with electric potentials).
Let $d\geq 3.$ Consider potentials of the form $V_{\text{el}}(x):=v_{\text{el}}(x)\mathbb{I}.$ Let $v_{\text{el}}\in L^{1}_{\text{loc}}(\mathbb{R}^{d};\mathbb{R})$ be such that $v_{\text{el}}\in C^{\infty}(\mathbb{R}^{d}\setminus\{0\})$ and that $v_{\text{el}}$ is bounded outside any ball containing the origin. Additionally, assume that there exist constants $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4}\geq 0$ satisfying
$\left(\frac{4d-6}{d-2}\right)\varepsilon_{1}+\left(\frac{4}{d-2}\right)\varepsilon_{2}\,\varepsilon_{3}+\left(\frac{8d-8}{d-2}\right)\varepsilon_{3}+2\varepsilon_{2}+2m\varepsilon_{4}^{2}<2,$ (1.14)
and assume (1.5)-(1.8). Then $\sigma_{p}(H_{m}(A,V_{\text{el}}))=\varnothing$.
The previous result has as immediate corollary the analogous result in the massless case, namely when $m=0$. Nevertheless, in this case, different manipulation of the eigenvalue equation gives a stronger result than Theorem 1.3.
Theorem 1.4 (Massless Dirac operator with electric potentials).
Let $d\geq 3.$ Consider potentials of the form $V_{\text{el}}(x):=v_{\text{el}}(x)\mathbb{I}.$ Let $v_{\text{el}}\in L^{1}_{\text{loc}}(\mathbb{R}^{d};\mathbb{R})$ be such that $v_{\text{el}}\in C^{\infty}(\mathbb{R}^{d}\setminus\{0\})$ and that $v_{\text{el}}$ is bounded outside any ball containing the origin. Additionally, assume that there exist constants $\varepsilon_{1},\varepsilon_{2}\geq 0$ satisfying
$\Big{(}\frac{4d-7}{d-2}\Big{)}\varepsilon_{1}+2\varepsilon_{2}+\varepsilon_{2}^{2}<1$
and assume (1.5) and (1.6). Then $\sigma_{p}(H_{0}(A,V_{\text{el}}))=\varnothing$.
Theorem 1.5 (Scalar potentials).
Let $d\geq 3.$ Consider potentials of the form $V_{\text{sc}}(x):=v_{\text{sc}}(x)\beta.$ Let $v_{\text{sc}}\in L^{1}_{\text{loc}}(\mathbb{R}^{d};\mathbb{R})$ be such that $v_{\text{sc}}\in C^{\infty}(\mathbb{R}^{d}\setminus\{0\})$ and that $v_{\text{sc}}$ is bounded outside any ball containing the origin. Additionally, assume that there exist constants $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4}\geq 0$ satisfying
$\left(\frac{4d-6}{d-2}\right)\varepsilon_{1}+2\varepsilon_{2}\,\varepsilon_{3}+2m\varepsilon_{4}^{2}+\left(\frac{4d-4}{d-2}\right)\varepsilon_{2}<2,$ (1.15)
and assume (1.5), (1.6) and, if $m\neq 0,$ assume also (1.8). Moreover assume that
$\int|v_{\text{sc}}|^{2}|\psi|^{2}\leq\varepsilon_{3}^{2}\|\nabla_{\!\text{A}}\psi\|^{2}$ (1.16)
holds true for every $\psi\in H^{1}(\mathbb{R}^{d})$. Then $\sigma_{p}(H_{m}(A,V_{\text{sc}}))=\varnothing$.
Theorem 1.6 (Anomalous magnetic potentials).
Let $d\geq 3.$ Consider potentials of the form $V_{\text{am}}(x):=i\beta(\boldsymbol{\alpha}\cdot\nabla)\phi_{\text{am}}(x).$ Let $\phi_{\text{am}}\in L^{2}_{\text{loc}}(\mathbb{R}^{d};\mathbb{R})$ be such that $\phi_{\text{am}}\in C^{\infty}(\mathbb{R}^{d}\setminus\{0\})$ and that $\phi_{\text{am}}$ is bounded outside any ball containing the origin. Additionally, assume that there exist constants $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\geq 0$ satisfying
$\left(\frac{4d-6}{d-2}\right)\varepsilon_{1}+\left(\frac{16d-16}{d-2}\right)\varepsilon_{2}+\left(\frac{8d-8}{(d-2)^{2}}\right)\varepsilon_{2}^{2}+\left(\frac{4d-4}{d-2}\right)\varepsilon_{3}<2,$ (1.17)
and assume (1.5). Moreover assume that
$\sup_{x\in\mathbb{R}^{d}}|x||\nabla\phi_{\text{am}}|$ $\leq\varepsilon_{2},$ (1.18)
$\int|x|^{2}|\Delta\phi_{\text{am}}|^{2}|\psi|^{2}$ $\leq\varepsilon_{3}^{2}\|\nabla_{\!\text{A}}\psi\|^{2},$ (1.19)
hold true for every $\psi\in H^{1}(\mathbb{R}^{d})$. Then $\sigma_{p}(H_{m}(A,V_{\text{am}}))=\varnothing$.
In three dimensions, we can consider an anomalous magnetic potential involving the spin operator, namely $V_{\text{am}}^{3d}:=i\beta(\boldsymbol{\alpha}\cdot\nabla)\phi_{\text{am}}-2\beta\boldsymbol{S}\cdot B,$ where $\boldsymbol{S}=(S_{1},S_{2},S_{3})$ represents the spin operator which is defined in terms of the Pauli matrices $\sigma_{j},$ $j=1,2,3$ as follows
$S_{j}:=\frac{1}{2}\begin{pmatrix}\sigma_{j}&0\\ 0&\sigma_{j}\end{pmatrix},\quad j=1,2,3.$
Theorem 1.7 (Anomalous magnetic, $d=3$).
Let $d=3.$ Consider potentials of the form $V_{\text{am}}^{3d}(x):=i\beta(\boldsymbol{\alpha}\cdot\nabla)\phi_{\text{am}}(x)-2\beta\boldsymbol{S}\cdot B(x).$ Let $\phi_{\text{am}}\in L^{2}_{\text{loc}}(\mathbb{R}^{3};\mathbb{R})$ be such that $\phi_{\text{am}}\in C^{\infty}(\mathbb{R}^{3}\setminus\{0\})$ and that $\phi_{\text{am}}$ is bounded outside any ball containing the origin. Additionally, assume that there exist constants $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4},\varepsilon_{5}\geq 0$ satisfying
$\Big{(}\frac{4d-6}{d-2}\Big{)}\varepsilon_{1}+\Big{(}\frac{8d-8}{(d-2)^{2}}\Big{)}\varepsilon_{2}^{2}+\Big{(}\frac{8d-8}{(d-2)^{2}}\Big{)}\varepsilon_{4}^{2}+\Big{(}\frac{16d-16}{(d-2)^{2}}\Big{)}\varepsilon_{2}\,\varepsilon_{4}+m\Big{(}\frac{16d-16}{d-2}\Big{)}\varepsilon_{1}\\ +\Big{(}\frac{16d-16}{d-2}\Big{)}\varepsilon_{2}+\Big{(}\frac{24d-24}{d-2}\Big{)}\varepsilon_{4}+\Big{(}\frac{4d-4}{d-2}\Big{)}\varepsilon_{3}+\Big{(}\frac{8d-8}{d-2}\Big{)}\varepsilon_{5}<2,$ (1.20)
and assume (1.5), (1.18) and (1.19). Moreover assume that
$\sup_{x\in\mathbb{R}^{d}}|x||B(x)|$ $\leq\varepsilon_{4},$ (1.21)
$\int|x|^{2}|\nabla B|^{2}|\psi|^{2}$ $\leq\varepsilon_{5}^{2}\|\nabla_{\!\text{A}}\psi\|^{2},$ (1.22)
hold true for every $\psi\in H^{1}(\mathbb{R}^{3})$. Then $\sigma_{p}(H_{m}(A,V_{\text{am}}^{3d}))=\varnothing$.
Remark 1.5 (Spectral stability).
The results presented above provide sufficient conditions for the potentials to ensure absence of eigenvalues of the perturbed Hamiltonian (1.1), which is crucial in some mathematical and quantum mechanical problems. In other words, we give sufficient conditions to ensure stability of the point spectrum under suitable perturbations.
In the purely electric case, namely if $A=0,$ one can actually prove stability of the entire spectrum. Indeed, condition (1.7) and Weyl’s theorem ensure that $\sigma_{\text{ess}}(H_{m}(0,V))=\sigma_{\text{ess}}(H_{m})=(-\infty,m]\cup[m,\infty).$ Moreover, since the residual spectrum is always empty in the self-adjoint case, we finally have $\sigma(H_{m}(0,V))=\sigma(H_{m}).$ In order to obtain a similar result in the electromagnetic case, it would be enough to assume that $\lim_{|x|\to\infty}|A(x)|=0.$ This does not, in principle, follow from our conditions on $A,$ but this property does turn out to be true in many interesting models.
1.1. Organization of the paper.
The paper is organized as follows: In Section 2 we provide a precise definition of the electromagnetic Dirac operator (1.1), giving also some minimal hypothesis on the potentials to ensure the self-adjointness of this operator in the appropriate domain. In Section 3 we provide some technical results needed to make rigorous all the computations. Finally, section 4 is dedicated to the proofs of the results presented in the introduction.
1.2. Notation.
Here we compile some notation and useful properties used in this work.
  • For a given magnetic potential $A:\mathbb{R}^{d}\to\mathbb{R}^{d}$ we define the magnetic gradient $\nabla_{\!\text{A}}=\nabla-iA$, which satisfies the following product rule
    $\nabla_{\!\text{A}}(fg)=g\nabla_{\!\text{A}}f+f\nabla g.$ (1.23)
    We also define the magnetic field associated with $A$ as the matrix-valued function given by
    $B_{jk}(x)=\partial_{j}A_{k}(x)-\partial_{k}A_{j}(x),$
    where the derivatives are taken in distributional sense. In dimension $d=3$ we can identify $B$ with the unique vector $\hat{B}\in\mathbb{R}^{3}$ such that
    $Bv=\hat{B}\times v$
    for all $v\in\mathbb{R}^{3}$, and in this case we are not going to make distinction between $B$ and $\hat{B}$.
    We refer to [5] for a complete survey on the concept of magnetic field in any dimensions and its definition in terms of differential forms and tensor fields.
    It is easy to see that the magnetic derivatives satisfy the following commutation relation
    $\partial_{j}^{A},\partial_{k}^{A}]=iB_{jk},\quad j,k\in\{1,2,\dots,d\}.$ (1.24)
    We will also use that
    $2\Re(\overline{\psi}\partial_{j}^{A}\psi)=2\Re(\overline{\psi}\partial_{j}\psi)=\partial_{j}(|\psi|^{2}).$ (1.25)
  • For a column vector $v\in\mathbb{C}^{d}$ its adjoint $v^{*}$ is the row vector given by
    $(v^{*})_{j}=\bar{v}_{j}.$
    With this, we have that the usual norm in $\mathbb{C}^{d}$ can be written as $|v|^{2}=v^{*}v$. For matrices we use the same symbol for its norm, that is, if $M$ is a $d\times d$ matrix we set
    $|M|=\sup_{v\in\mathbb{C}^{d}\setminus\{0\}}\frac{|Mv|}{|v|}.$
  • For the sake of simplicity, we will write $L^{2}(\mathbb{R}^{d})$ instead of $L^{2}(\mathbb{R}^{d};\mathbb{C}^{N})$ for vector valued functions. In the same spirit, the symbols $\|\cdot\|$ and $\langle\cdot,\cdot\rangle$ are used for the norm and scalar product in $L^{2}(\mathbb{R}^{d}),$ respectively.
  • For two vectors $u,v$ we use the centered dot to denote the Euclidean scalar product, that is
    $v\cdot u=\sum_{j=1}^{d}v_{j}u_{j}.$
    This notation will be used regardless of the nature of $v_{j}$ and $u_{j}$. Thus, for example, if $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{d})$ is a vector of matrices, we write
    $\begin{split}\boldsymbol{\alpha}\cdot\nabla&=\sum_{j=1}^{d}\alpha_{j}\partial_{j},\\ \boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha}&=\sum_{j,k=1}^{d}B_{jk}\alpha_{j}\alpha_{k}.\end{split}$
2. Definition of Dirac operator
In this section we gather some basic properties on the Dirac operators. We refer to [36] for a more exhaustive dissertation.
Following the notation used in the introduction, the free Dirac operator in dimension $d\geq 3$ is defined as
$H_{m}\coloneqq H_{m}(0,0)=-i\boldsymbol{\alpha}\cdot\nabla+m\beta,$ (2.1)
where $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{d})$ and $\beta$ are the so-called Dirac matrices. These are $N\times N$ Hermitian matrices, with $N=2^{\lfloor(d+1)/2\rfloor}$, which satisfy the following anti-commutation relations
$\begin{split}\alpha_{j}\alpha_{k}+\alpha_{k}\alpha_{j}&=2\delta_{jk}\mathbb{I}_{N},\\ \alpha_{j}\beta+\beta\alpha_{j}&=0,\\ \beta^{2}&=\mathbb{I}_{N},\end{split}$
for $j,k\in\{1,2,\dots,d\}$. For an explicit construction of these matrices in any dimension see for example [6]. It is well known that this operator is self-adjoint in $L^{2}(\mathbb{R}^{d})$ with domain $D(H_{m})=H^{1}(\mathbb{R}^{d})$ and has pure continuous spectrum, namely
$\sigma(H_{m})=\sigma_{c}(H_{m})=(-\infty,-m]\cup[m,\infty),$
see for example [36, Sec.1.4.4].
In this work we are interested in perturbing (2.1) with a magnetic potential $A:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and an electric potential $V:\mathbb{R}^{d}\to\mathbb{C}^{N\times N}.$ More precisely, we are interested in the electromagnetic Dirac operator defined in (1.1).
In order to have that (1.1) is well defined as a self-adjoint operator in $L^{2}(\mathbb{R}^{d}),$ from now on, we will assume that $A\in L^{2}_{loc}(\mathbb{R}^{d})$ is such that $B\in L^{2}_{loc}(\mathbb{R}^{d})$ and such that $H_{m}(A):=H_{m}(A,0)$ is self-adjoint in $L^{2}(\mathbb{R}^{d})$ with domain $H^{1}(\mathbb{R}^{d})$. This class of potentials includes, for example, any smooth magnetic potential, or vector potentials that satisfy
$|A(x)|\leq\frac{a}{|x|}+b,$
for $a>0$ small enough (see [36, Sec.4.3]). For the electric potential we shall assume that $V(x)$ is Hermitian for almost all $x\in\mathbb{R}^{d}$. Furthermore, to ensure the self-adjointness of $H_{m}(A,V)$ and, in consequence, that its spectrum is real, we will assume that $V$ is $H_{m}(A)$-bounded with relative bound less than 1. This means that there exist constants $a,b>0$ with $a\leq 1$ such that for every $\psi\in H^{1}(\mathbb{R}^{d})$ the following inequality
$\|V\psi\|^{2}\leq a\|H_{m}(A)\psi\|^{2}+b\|\psi\|^{2},$ (2.2)
holds. This allows us to define $H_{m}(A,V)$ as a self-adjoint operator in $L^{2}(\mathbb{R}^{d})$ with domain $H^{1}(\mathbb{R}^{d})$ by the Kato-Rellich theorem [29, Sec.X.2]. Notice that in some specific cases of special interest, condition (2.2) can be replaced by weaker assumptions.
In the next result we show that condition (2.2) implies that $V$ is $\nabla_{\!\text{A}}$-bounded.
Lemma 2.1.
Let $V\colon\mathbb{R}^{d}\to\mathbb{C}^{N\times N}$ and $A\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ satisfy (2.2) and (1.5), respectively. Then there exist constants $\mu,\nu>0$ such that
$\|V\psi\|^{2}\leq\mu\|\nabla_{\!\text{A}}\psi\|^{2}+\nu\|\psi\|^{2}$ (2.3)
holds for all $\psi\in H^{1}(\mathbb{R}^{d})$.
Proof.
By density, we consider $\psi\in C^{\infty}_{0}(\mathbb{R}^{d}).$ First of all we observe that
$\begin{split}\|H_{m}(A)\psi\|^{2}&=\langle\psi,H_{m}(A)^{2}\psi\rangle\\ &=\langle\psi,-\Delta_{\text{A}}\psi\rangle+\langle\psi,-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi\rangle+m^{2}\|\psi\|^{2}\\ &=\|\nabla_{\!\text{A}}\psi\|^{2}+\langle\psi,-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi\rangle+m^{2}\|\psi\|^{2}.\end{split}$
Using (1.5) and the Hardy inequality (1.9) one has
$\begin{split}\langle\psi,-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi\rangle&\leq\frac{1}{2}\int|B||\psi|^{2}\\ &\leq\frac{1}{2}\left(\int|x|^{2}|B|^{2}|\psi|^{2}\right)^{1/2}\left(\int\frac{|\psi|^{2}}{|x|^{2}}\right)^{1/2}\\ &\leq\frac{\varepsilon_{1}}{d-2}\|\nabla_{\!\text{A}}\psi\|^{2}.\end{split}$
Plugging the latter inequality in the previous identity, we have
$\|H_{m}(A)\psi\|^{2}\leq\left(1+\frac{\varepsilon_{1}}{d-2}\right)\|\nabla_{\!\text{A}}\psi\|^{2}+m^{2}\|\psi\|^{2}.$ (2.4)
Inserting this into (2.2) we obtain (2.3) with $\mu=a\big{(}1+\frac{\varepsilon_{1}}{d-2}\big{)}$ and $\nu=am^{2}+b.$
3. Preliminary results
As mentioned in the introduction, the main ingredient in the proof of Theorem 1.2 is the method of multipliers. The idea behind this method is to construct suitable weighted identities for solutions to the eigenvalue equation and then use smallness assumptions on the potentials to arrive to a contradiction. To make the argument rigorous, we will need some preliminary lemmas contained in this section.
As a first step we need to approximate any solution $\psi$ of the eigenvalue equation with a sequence of compactly supported functions $\psi_{R}$ with compact support in $\mathbb{R}^{d}\setminus\{0\}.$ To achieve this, we define $\xi\in C_{0}^{\infty}(\mathbb{R}^{d})$ such that
$\xi(x)=\begin{cases}1&\text{ if }|x|\leq 1,\\ 0&\text{ if }|x|\geq 2,\end{cases}$
and for $R>0$, we define
$\xi_{R}(x)=\xi\left(\frac{x}{R}\right)-\xi(Rx).$ (3.1)
Then we have that
$\nabla\xi_{R}(x)=\frac{1}{R}\nabla\xi\left(\frac{x}{R}\right)-R\nabla\xi(Rx),\quad\text{and}\quad\Delta\xi_{R}(x)=\frac{1}{R^{2}}\Delta\xi\left(\frac{x}{R}\right)-R^{2}\Delta\xi(Rx).$
Notice that both $\nabla\xi(Rx)$ and $\Delta\xi(Rx)$ are supported in the anulus
$D(R)=\left\{x\in\mathbb{R}^{d}\colon\frac{1}{R}\leq|x|\leq\frac{2}{R}\right\},$ (3.2)
and, similarly, $\nabla\xi\left(\frac{x}{R}\right)$ and $\Delta\xi\left(\frac{x}{R}\right)$ are supported in $B(2R)$, the ball centered at 0 with radius $2R.$
For any spinorial function $f:\mathbb{R}^{d}\to\mathbb{C}^{N}$ we define the compactly supported approximating family of functions with compact support in $\mathbb{R}^{d}\setminus\{0\}$ as $f_{R}=\xi_{R}f$.
If $\psi\in H^{1}(\mathbb{R}^{d})$ is a solution of
$i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi+m\beta\psi+V\psi=\lambda\psi,$ (3.3)
than it is easy to show that $\psi_{R}\in H^{1}(\mathbb{R}^{d})$ and solves the approximating equation
$i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi_{R}+m\beta\psi_{R}+V\psi_{R}=\lambda\psi_{R}-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi.$ (3.4)
The strategy to prove Theorem 1.2 is to reduce to prove an analogous result for solutions to a second order eigenvalues equation. In the next result we will show that if $\psi_{R}\in H^{1}(\mathbb{R}^{d})$ is a solution to the first order equation (3.4), then it is more regular and therefore satisfies a second-order associated problem.
Lemma 3.1.
Let $A$ and $V$ satisfy the hypothesis of Theorem 1.2. If $\psi_{R}\in H^{1}(\mathbb{R}^{d})$ is a solution of (3.4), then $\psi_{R}\in H^{2}(\mathbb{R}^{d})$ and satisfies
$H_{m}(A)^{2}\psi_{R}+V^{2}\psi_{R}+m\{\beta,V\}\psi_{R}-i(\{\boldsymbol{\alpha},V\}\cdot\nabla_{\!\text{A}})\psi_{R}-i(\boldsymbol{\alpha}\cdot\nabla V)\psi_{R}\\ =\lambda^{2}\psi_{R}+(H_{m}(A,V)+\lambda)(-i(\boldsymbol{\alpha}\cdot\nabla)\xi_{R})\psi.$ (3.5)
Proof.
Notice that in order to prove $\psi_{R}\in H^{2}(\mathbb{R}^{d})$ it is enough to show that $(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}})\psi_{R}\in H^{1}(\mathbb{R}^{d})$. Since $\psi_{R}$ satisfies (3.4) then one needs to show that $V\psi_{R}\in H^{1}(\mathbb{R}^{d})$ since the other terms in the identity are trivially in $H^{1}(\mathbb{R}^{d}).$ Observe that since $\text{supp}(\xi_{R})\subseteq\{x\in\mathbb{R}^{d}:R^{-1}\leq|x|\leq 2R\},$ the same holds for $\psi_{R}$ by definition. Since $V$ is smooth and bounded in this set then $V\psi_{R}\in H^{1}(\mathbb{R}^{d}).$
To prove (3.5), we apply $H_{m}(A,V)$ to both sides of (3.4). This is a rigorous step since we have just shown that all terms in the equation belong to $H^{1}(\mathbb{R}^{d})$. We consider first the left hand side.
$H_{m}(A,V)^{2}\psi_{R}=[H_{m}(A)^{2}+V^{2}+\{H_{m}(A),V\}]\psi_{R}.$
Using the bi-linearity of the anti-commutator and the product rule for the magnetic gradient (1.23) one has
$\{H_{m}(A),V\}=\{-i(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}),V\}+m\{\beta,V\}=-i(\{\boldsymbol{\alpha},V\}\cdot\nabla_{\!\text{A}})-i(\boldsymbol{\alpha}\cdot\nabla)V+m\{\beta,V\}.$
For the right hand side we have
$\begin{split}\lambda H_{m}(A,V)\psi_{R}+H_{m}(A,V)(-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi)&=\lambda(\lambda\psi_{R}-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi)+H_{m}(A,V)(-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi)\\ &=\lambda^{2}\psi_{R}+(H_{m}(A,V)+\lambda)(-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi),\end{split}$
where we have simply used that $\psi_{R}$ satisfies (3.4). Gathering the two identities gives the proof. ∎
The last term in the right-hand-side of (3.5) can be seen as an error term that comes from approximating $\psi$ by $\psi_{R}$. The following proposition shows that this error term becomes negligible as $R$ goes to infinity in a suitable sense dictated by the manipulation that we will make to (3.5).
Proposition 3.2.
Suppose that $V\colon\mathbb{R}^{d}\to\mathbb{C}^{N\times N}$ is such that $V\in\mathbb{C}^{\infty}(\mathbb{R}^{d}\setminus\{0\})$ and that $V$ is bounded outside any ball containing the origin. Moreover, assume (2.2). If $\psi\in H^{1}(\mathbb{R}^{d})$ solves (3.3) then
$\lim_{R\to\infty}\langle 2(x\cdot\nabla_{\!\text{A}})\psi_{R}+d\psi_{R},(H_{m}(A,V)+\lambda)(-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi)\rangle=0.$ (3.6)
In order to prove Proposition 3.2 we will need to show that the $L^{2}$ norm of $(\nabla\xi_{R})\psi$ becomes negligible as $R$ goes to $\infty.$ We will also need to show that $\psi_{R}$ converges to $\psi$ in $L^{2}(\mathbb{R}^{d})$ and that $\nabla_{\!\text{A}}\psi_{R}$ converges to $\nabla\psi.$ This is proved in the next lemma.
Lemma 3.3.
Let $\xi_{R}$ be as in (3.1) above. Then for any $\psi\in H^{1}(\mathbb{R}^{d})$ we have
$\lim_{R\to\infty}\|(\nabla\xi_{R})\psi\|=0,$ (3.7)
$\lim_{R\to\infty}\|\nabla_{\!\text{A}}\psi-\nabla_{\!\text{A}}\psi_{R}\|+\|\psi-\psi_{R}\|=0.$ (3.8)
Proof.
Let us see (3.7) first. By definition, we have
$(\nabla\xi_{R})\psi=\frac{1}{R}\nabla\xi\left(\frac{x}{R}\right)\psi-R\nabla\xi(Rx)\psi.$
Since $\nabla\xi$ is bounded, the first term converges to 0 in $L^{2}(\mathbb{R}^{d})$. For the other term we use that $\nabla\xi(Rx)$ is supported in the anulus $D(R)$ defined in (3.2), and hence
$\begin{split}\|R\nabla\xi(Rx)\psi\|^{2}&=\int_{D(R)}R^{2}|\nabla\xi(Rx)|^{2}|\psi|^{2}\\ &\leq 4\int_{D(R)}|\nabla\xi(Rx)|^{2}\frac{|\psi|^{2}}{|x|^{2}}\\ &\leq 4\|\nabla\xi\|_{\infty}^{2}\int_{D(R)}\frac{|\psi|^{2}}{|x|^{2}}.\end{split}$
Therefore, we only need to prove that the last integral converges to 0. From the Hardy inequality we have that $\frac{|\psi|^{2}}{|x|^{2}}$ is integrable, and since the measure of $D(R)$ converges to 0 as $R$ goes to infinity, we obtain the desired limit.
To prove (3.8), observe that from the definition of $\xi_{R}$ it follows that $\xi_{R}\to 1$ a.e. and $|\xi_{R}|\leq 2.$ Therefore, $\psi_{R}\to\psi$ in $L^{2}(\mathbb{R}^{d})$ by the dominated convergence theorem. In order to show that $\nabla_{\!\text{A}}\psi_{R}\to\nabla_{\!\text{A}}\psi$ in $L^{2}(\mathbb{R}^{d})$ one observes that
$\nabla_{\!\text{A}}\psi_{R}=\xi_{R}\nabla_{\!\text{A}}\psi+(\nabla\xi_{R})\psi.$
Now the first term converges to $\nabla_{\!\text{A}}\psi$ in $L^{2}(\mathbb{R}^{d})$ using the same argument as above. For the second term, we have just shown that it converges to zero. This concludes (3.8) and the proof of the lemma. ∎
The proof of Proposition 3.2 immediately follows from the following lemma.
Lemma 3.4.
Under the hypotheses of proposition 3.2 one has
$\lim_{R\to\infty}\langle\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle=0,$ (3.9)
$\lim_{R\to\infty}\langle\psi_{R},H_{m}(A,V)(-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi)\rangle=0,$ (3.10)
$\lim_{R\to\infty}\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R},H_{m}(A)(-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi)\rangle=0,$ (3.11)
$\lim_{R\to\infty}\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle=0,$ (3.12)
$\lim_{R\to\infty}\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R},V((-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi)\rangle=0.$ (3.13)
Proof.
The first limit (3.9) simply follows from the Cauchy-Schwarz inequality and from (3.7) in Lemma 3.3. For the second identity (3.10) we observe
$\begin{split}\langle\psi_{R},H_{m}(A,V)(-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi)\rangle=&\langle H_{m}(A)\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle,+\langle V\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle\\ \leq&\|H_{m}(A)\psi_{R}\|\|(\nabla\xi_{R})\psi\|+\|V\psi_{R}\|\|\|(\nabla\xi_{R})\psi\|.\end{split}$
From estimates (2.3) and (2.4) in Lemma 2.1, we have that both $\|H_{m}(A)\psi_{R}\|$ and $\|V\psi_{R}\|$ are uniformly bounded in $R$ because, from Lemma 3.3, $\psi_{R}$ converges in $H^{1}_{A}(\mathbb{R}^{d})$. This allows us to conclude (3.10) since $(\nabla\xi_{R})\psi$ converges to 0 in $L^{2}(\mathbb{R}^{d})$ as it is shown in (3.7) of Lemma 3.3.
We continue proving (3.11). One has
$H_{m}(A)(-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi)=-\Delta\xi_{R}\psi-\alpha_{j}\alpha_{k}\partial_{k}\xi_{R}\partial_{j}^{A}\psi-im\beta(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi,$
and then we have
$\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R},H_{m}(A)(-i(\boldsymbol{\alpha}\cdot\nabla)\xi_{R}\psi)\rangle\\ =\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R},-\Delta\xi_{R}\psi\rangle+\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R},-\alpha_{j}\alpha_{k}\partial_{k}\xi_{R}\partial_{j}^{A}\psi\rangle+\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R},-im\beta(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle.$ (3.14)
We shall estimate each term of the right-hand-side of (3.14) separately. First,
$\begin{split}\left|\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R},-\Delta\xi_{R}\psi\rangle\right|&\leq 2\int|x||\nabla_{\!\text{A}}\psi_{R}||\psi||\Delta\xi_{R}|\\ &\leq 2\int_{D(R)}R^{2}|x||\nabla_{\!\text{A}}\psi_{R}||\psi||\Delta\xi(Rx)|+\frac{2}{R}\int_{B(2R)}\frac{|x|}{R}|\nabla_{\!\text{A}}\psi_{R}||\psi|\left|\Delta\xi\left(\frac{x}{R}\right)\right|\\ &\leq 8\|\Delta\xi\|_{\infty}\int_{D(R)}|\nabla_{\!\text{A}}\psi_{R}|\frac{|\psi|}{|x|}+\frac{4\|\Delta\xi\|_{\infty}}{R}\int_{B(2R)}|\nabla_{\!\text{A}}\psi_{R}||\psi|.\end{split}$
From Lemma 3.3 we know that $\|\nabla_{\!\text{A}}\psi_{R}\|$ is bounded. Moreover, since $\tfrac{|\psi|}{|x|}$ is square integrable, the first integral converges to 0 because the measure of $D(R)$ does. The convergence to zero of the second integral is immediate.
Now we shall see the second term in the right-hand-side of (3.14). Similarly as above one has
$\begin{split}|\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R},-\alpha_{j}\alpha_{k}\partial_{k}\xi_{R}\partial_{j}^{A}\psi\rangle|&\leq 2\int|x||\nabla_{\!\text{A}}\psi_{R}||\nabla_{\!\text{A}}\psi||\nabla\xi_{R}|\\ &\leq 2\int_{D(R)}R|x||\nabla_{\!\text{A}}\psi_{R}||\nabla_{\!\text{A}}\psi|\left|\nabla\xi\left(Rx\right)\right|+2\int_{B(2R)}\frac{|x|}{R}|\nabla_{\!\text{A}}\psi_{R}||\nabla_{\!\text{A}}\psi|\left|\nabla\xi\left(\frac{x}{R}\right)\right|\\ &\leq 4\|\nabla\xi\|_{\infty}\int_{D(R)}|\nabla_{\!\text{A}}\psi_{R}||\nabla_{\!\text{A}}\psi|+4\int_{B(2R)}|\nabla_{\!\text{A}}\psi_{R}||\nabla_{\!\text{A}}\psi|\left|\nabla\xi\left(\frac{x}{R}\right)\right|.\end{split}$
The first integral converges to 0 in the same way as above. For the second we use that $\nabla\xi(\tfrac{x}{R})$ converges to 0 and is uniformly bounded in $R$. We conclude by the dominated convergence theorem.
For the last term in the right hand side of (3.14), we have
$\begin{split}|\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R},-im\beta(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle|&\leq 2m\int|x||\nabla_{\!\text{A}}\psi_{R}||\psi||\nabla\xi_{R}|\\ &\leq 2m\int_{D(R)}R|x||\nabla_{\!\text{A}}\psi_{R}||\psi||\nabla\xi(Rx)|+2m\int_{B(2R)}\frac{|x|}{R}|\nabla_{\!\text{A}}\psi_{R}||\psi|\left|\nabla\xi\left(\frac{x}{R}\right)\right|\\ &\leq 4m\|\nabla\xi\|_{\infty}\int_{D(R)}|\nabla_{\!\text{A}}\psi_{R}||\psi|+4m\int_{B(2R)}|\nabla_{\!\text{A}}\psi_{R}||\psi|\left|\nabla\xi\left(\frac{x}{R}\right)\right|.\end{split}$
Hence, by using similar arguments as before, we conclude the proof of (3.11).
Notice that the last estimate also provides the proof of (3.12).
Finally, to prove (3.13) one has
$\begin{split}|\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R},V((-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi)\rangle|&\leq 2\int|x||\nabla_{\!\text{A}}\psi_{R}||V||\psi||\nabla\xi_{R}|\\ &\leq 2\int_{D(R)}R|x||\nabla_{\!\text{A}}\psi_{R}||V||\psi||\nabla\xi(Rx)|+2\int_{B(2R)}\frac{|x|}{R}|\nabla_{\!\text{A}}\psi_{R}||V||\psi|\left|\nabla\xi\left(\frac{x}{R}\right)\right|\\ &\leq 4\|\nabla\xi\|_{\infty}\int_{D(R)}|\nabla_{\!\text{A}}\psi_{R}||V||\psi|+4\int_{B(2R)}|\nabla_{\!\text{A}}\psi_{R}||V||\psi|\left|\nabla\xi\left(\frac{x}{R}\right)\right|,\end{split}$
where the convergence to zero in ensured reasoning as before using that $V$ satisfies (2.3). This concludes the proof of the lemma. ∎
Proof of Proposition 3.2.
The proof of (3.6) is an immediate consequence of the validity of (3.9)-(3.13) in Lemma 3.4. ∎
Lemma 3.5.
Let $M\in W^{1,1}_{loc}(\mathbb{R}^{d},\mathbb{R}^{N\times N})$ be a matrix-valued function which is Hermitian almost everywhere. Moreover assume $\psi\in H^{2}(\mathbb{R}^{d})$ to be compactly supported and $c\in\mathbb{R}$. Then the following identities hold
$\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi+d\psi,-\Delta_{A}\psi\rangle$ $=2\|\nabla_{\!\text{A}}\psi\|^{2}+2\Im\int x_{k}B_{jk}\psi^{*}\partial_{j}^{A}\psi,$ (3.15)
$\Re\langle 2x\cdot\nabla\psi+d\psi,M\psi\rangle$ $=-\langle\psi,(x\cdot\nabla M)\psi\rangle,$ (3.16)
$\Re\langle 2x\cdot\nabla\psi+d\psi,c\psi\rangle$ $=0.$ (3.17)
Proof.
We start by proving (3.15). Integrating by parts, we obtain
$\begin{split}\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi+d\psi,-\Delta_{\text{A}}\psi\rangle&=\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi,-\Delta_{\text{A}}\psi\rangle-d\Re\langle\psi,-\Delta_{\text{A}}\psi\rangle\\ &=\Re\langle\nabla_{\!\text{A}}(2x\cdot\nabla_{\!\text{A}}\psi),\nabla_{\!\text{A}}\psi\rangle+d\|\nabla_{\!\text{A}}\psi\|^{2}.\end{split}$
By using the commutation relation (1.24) one easily gets
$\begin{split}\Re\langle\nabla_{\!\text{A}}(2x\cdot\nabla_{\!\text{A}}\psi),\nabla_{\!\text{A}}\psi\rangle&=\Re\langle\partial_{j}^{A}(2x_{k}\partial_{k}^{A}\psi),\partial_{j}^{A}\psi\rangle=\Re\langle 2x_{k}\partial_{k}^{A}\partial_{j}^{A}\psi+2ix_{k}B_{jk}\psi+2\partial_{j}^{A}\psi,\partial_{j}^{A}\psi\rangle\\ &=(-d+2)\|\nabla_{\!\text{A}}\psi\|^{2}+2\Im\int x_{k}B_{jk}\psi^{*}\partial_{j}^{A}\psi,\end{split}$
where we have used that $2\Re(\partial_{k}^{A}\partial_{j}^{A}\psi)^{*}\partial_{j}^{A}\psi=\partial_{k}^{A}|\nabla_{\!\text{A}}\psi|^{2}.$ Notice that summation over repeated indices is implicit. Plugging the latter in the former gives (3.15).
Let us prove now (3.16). First, we compute $\Re\langle 2x\cdot\nabla\psi,M\psi\rangle.$ Integrating by parts and using that $M$ is Hermitian, one has
$\begin{split}\Re\langle 2x\cdot\nabla\psi,M\psi\rangle&=-2d\langle\psi,M\psi\rangle-2\Re\langle x_{j}\psi,\partial_{j}M\psi\rangle-2\Re\langle x_{j}\psi,M\partial_{j}\psi\rangle\\ &=-2d\langle\psi,M\psi\rangle-2\Re\langle\psi,(x\cdot\nabla M)\psi\rangle-\Re\langle M\psi,2x\cdot\nabla\psi\rangle.\end{split}$
This gives
$\Re\langle 2x\cdot\nabla\psi,M\psi\rangle=-d\langle\psi,M\psi\rangle-\Re\langle\psi,(x\cdot\nabla M)\psi\rangle.$
And, identity (3.16) follows immediately. We finally turn to (3.17), which trivially follows from (3.16) observing that any constant $c\in\mathbb{R}$ belongs to $W^{1,1}_{loc}(\mathbb{R}^{d}).$
4. Proof of main theorems
Proof of Theorem 1.1.
By contradiction suppose that there exist $\lambda\in(-m,m)$ and $\psi\in H^{1}(\mathbb{R}^{d})$ such that
$H_{m}(A,V)\psi=\lambda\psi.$
In particular, it follows that
$\|H_{m}(A,V)\psi\|^{2}=\lambda^{2}\|\psi\|^{2}.$ (4.1)
Computing the left-hand side gives
$\|H_{m}(A,V)\psi\|^{2}=\|H_{m}(A)\psi\|^{2}+\|V\psi\|^{2}+2\Re\langle V\psi,H_{m}(A)\psi\rangle,$
moreover
$\begin{split}\|H_{m}(A)\psi\|^{2}&=\|-i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi\|^{2}+m^{2}\|\psi\|^{2}+2\Re\langle-i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi,m\beta\psi\rangle\\ &=\|-i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi\|^{2}+m^{2}\|\psi\|^{2},\end{split}$
where in the last identity we have used that $\Re\langle-i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi,m\beta\psi\rangle=0$ since $i(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}})$, and $\beta$ are anti-commuting self-adjoint operators.
Plugging the latter in the former gives
$\|H_{m}(A,V)\psi\|^{2}=\|-i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi\|^{2}+m^{2}\|\psi\|^{2}+\|V\psi\|^{2}+2\Re\langle V\psi,H_{m}(A)\psi\rangle.$
Using this in (4.1) we have
$\|-i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi\|^{2}+\|V\psi\|^{2}+2\Re\langle V\psi,H_{m}(A)\psi\rangle=(\lambda^{2}-m^{2})\|\psi\|^{2}.$
From the Cauchy-Schwartz inequality, we see that the left hand side of the equation above is bounded from below by
$\|-i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi\|^{2}+\|V\psi\|^{2}-2\|V\psi\|\|H_{m}(A)\psi\|$ $=\left(\|H_{m}(A)\psi\|-\|V\psi\|\right)^{2}+\|-i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi\|^{2}-\|H_{m}(A)\psi\|^{2}$
$=(\|H_{m}(A)\psi\|-\|V\psi\|)^{2}-m^{2}\|\psi\|^{2}.$
The last term is positive by the assumption (1.3) on $V$. This implies
$(\lambda^{2}-m^{2})\|\psi\|^{2}\geq 0.$
Since $\lambda\in(-m,m),$ we have $\lambda^{2}-m^{2}<0.$ Therefore, the previous inequality implies that $\psi=0,$ which is a contradiction. ∎
Proof of Theorem 1.2.
Assume, by contradiction, that $\psi\in H^{1}(\mathbb{R}^{d})$ is a non trivial solution of the eigenvalue problem (3.3). From Lemma 3.1 we have that the approximated solution $\psi_{R}:=\xi_{R}\psi$ satisfies (3.5). Multiplying (3.5) by $2(x\cdot\nabla_{\!\text{A}})\psi_{R}+d\psi_{R},$ integrating over $\mathbb{R}^{d}$ and taking the real part of the corresponding identity, we obtain
$2\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+2\Im\int x_{k}B_{jk}\psi_{R}^{*}\partial_{j}^{A}\psi_{R}+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi_{R}\rangle\\ -\langle\psi_{R},(x\cdot\nabla)(V^{2})\psi_{R}\rangle-m\langle\psi_{R},(x\cdot\nabla\{\beta,V\})\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\{\boldsymbol{\alpha},V\}\cdot\nabla_{\!\text{A}})\psi_{R}\rangle+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla V)\psi_{R}\rangle\\ =\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},(H_{m}(A,V)+\lambda)-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle.$ (4.2)
Here, we have used Lemma 3.5 and that $V^{2}$ and $\{\beta,V\}$ are Hermitian. We start estimating the left hand side of (4.2). We consider first the terms depending on $B.$ Using the Cauchy-Schwarz inequality and assumption (1.5) one has
$2\left|\Im\int x_{k}B_{jk}\psi_{R}^{*}\partial_{j}^{A}\psi_{R}\right|\leq 2\int|x||B||\psi_{R}||\nabla_{\!\text{A}}\psi_{R}|\leq 2\varepsilon_{1}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}.$
In a similar way, using that the Dirac matrices are unitary and the standard Hardy inequality, we have
[…]
Similarly, for the terms involving the electric field $V,$ using (1.6) and (1.7), one has
$\begin{split}\left|\langle\psi_{R},(x\cdot\nabla)(V^{2})\psi_{R}\rangle\right|&\leq 2\int|x||V||\nabla V||\psi_{R}|^{2}\\ &\leq 2\left(\int|x|^{2}|\nabla V|^{2}|\psi_{R}|^{2}\right)^{1/2}\left(\int|V|^{2}|\psi_{R}|^{2}\right)^{1/2}\\ &\leq\left(\frac{4}{d-2}\right)\varepsilon_{2}\,\varepsilon_{3}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}.\end{split}$
For the commutators of two matrices $U,V$, we will use that $|\{U,V\}|\leq 2|U||V|.$ Using this and (1.8) one has
$\left|m\langle\psi_{R},(x\cdot\nabla\{\beta,V\})\psi_{R}\rangle\right|\leq 2m\int|x||\nabla V||\psi_{R}|^{2}\leq 2m\varepsilon_{4}^{2}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}.$
Moreover, using again (1.7), we have
$\begin{split}|\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\{\boldsymbol{\alpha},V\}\cdot\nabla_{\!\text{A}})\psi_{R}\rangle|&\leq 4\int|x||V||\nabla_{\!\text{A}}\psi_{R}|^{2}+2d\int|V||\psi_{R}||\nabla_{\!\text{A}}\psi_{R}|\\ &\leq 4\varepsilon_{3}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+\left(\frac{4d}{d-2}\right)\varepsilon_{3}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}\\ &\leq\left(\frac{8d-8}{d-2}\right)\varepsilon_{3}\|\nabla_{\!\text{A}}\psi_{R}\|^{2},\end{split}$
where we have used (1.6). We conclude estimating the last term of the left-hand-side of (4.2)
$\begin{split}|\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla V)\psi_{R}\rangle|&\leq 2\int|x||\nabla V||\psi_{R}||\nabla_{\!\text{A}}\psi_{R}|+d\int|\nabla V||\psi_{R}|^{2}\\ &\leq 2\varepsilon_{2}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+\left(\frac{2d}{d-2}\right)\varepsilon_{2}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}\\ &\leq\left(\frac{4d-4}{d-2}\right)\varepsilon_{2}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}.\end{split}$
Gathering all the previous estimates and inserting them in (4.2) we get
$\left(2-\left(\frac{4d-6}{d-2}\right)\varepsilon_{1}-\left(\frac{4}{d-2}\right)\varepsilon_{2}\,\varepsilon_{3}-2m\varepsilon_{4}^{2}-\left(\frac{8d-8}{d-2}\right)\varepsilon_{3}-\left(\frac{4d-4}{d-2}\right)\varepsilon_{2}\right)\|\nabla_{\!\text{A}}\psi_{R}\|^{2}\\ \leq\Re\langle 2(x\cdot\nabla_{\!\text{A}})\psi_{R}+d\psi_{R},(H_{m}(A,V)+\lambda)-i(\boldsymbol{\alpha}\cdot\nabla)\xi_{R}\psi\rangle$ (4.3)
Using (3.6) in Proposition 3.2 we have that the right hand side of (4.3) converges to 0, and in virtue of (1.4) we must have that $\psi$ is identically zero, which leads to a contradiction. ∎
Proof of Corollary 1.1.
Theorem 1.2 remains valid if we move the singularity of the potentials from 0 to another point in $\mathbb{R}^{d}$. To prove this, observe that if $V$ is such that $V\in C^{\infty}(\mathbb{R}^{d}\setminus\{x_{0}\})$ and $\psi\in H^{1}(\mathbb{R}^{d})$ is a solution of
$i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi+m\beta\psi+V\psi=\lambda\psi,$
then $\varphi(x)=\psi(x-x_{0})$ is a solution of
$i\boldsymbol{\alpha}\cdot\nabla_{\tilde{A}}\varphi++m\beta\varphi+\tilde{V}\varphi=\lambda\varphi,$
where $\tilde{A}(x)=A(x-x_{0})$ and $\tilde{V}(x)=V(x-x_{0})$. If $A$ and $V$ satisfy the hypotheses of the theorem, it is easy to check that $\tilde{A}$ and $\tilde{V}$ do too. ∎
Proof of Corollary 1.2.
Note also that if we prove that the Hamiltonian $i\alpha\cdot\nabla_{\!\text{A}}+m\beta+V$ has no eigenvalues, the same is true for any conjugation of $V$ by an unitary matrix. Indeed, if $\psi\in H^{1}(\mathbb{R}^{d})$ is solution of
$i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi+m\beta\psi+V\psi=\lambda\psi$
and $P$ is an unitary matrix, then $\varphi=P\psi$ is solution of
$i(P\boldsymbol{\alpha}P^{*})\cdot\nabla_{\!\text{A}}\varphi+m(P\beta P^{*})\varphi+(PVP^{*})\varphi=\lambda\varphi.$
Since conjugating the Dirac matrices produces another representation of them, the conjugated potential generates eigenvalues if and only if the original potential does. ∎
Proof of Theorem 1.3.
Writing the identity (4.2) in the proof of Theorem 1.2 for the specific case $V=V_{\text{el}}=v_{\text{el}}\mathbb{I}$, one has
$2\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+2\Im\int x_{k}B_{jk}\psi_{R}^{*}\partial_{j}^{A}\psi_{R}+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi_{R}\rangle\\ -\int x\cdot\nabla(v_{\text{el}}^{2})|\psi_{R}|^{2}-2m\langle\psi_{R},(x\cdot\nabla v_{\text{el}})\beta\psi_{R}\rangle+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},2v_{\text{el}}(-i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}})\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla v_{\text{el}})\psi_{R}\rangle=\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},(H_{m}(A,V_{\text{el}})+\lambda)-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle.$ (4.4)
One can easily convince themselves that the only term that can be estimated differently from the general case is the following
$\begin{split}I:&=\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla v_{\text{el}})\psi_{R}\rangle\\ &=\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla v_{\text{el}})\psi_{R}\rangle.\end{split}$ (4.5)
In the last identity we have used that since $i\boldsymbol{\alpha}\cdot\nabla v_{\text{el}}$ is antysimmetric, $\Re\langle\psi_{R},-i\boldsymbol{\alpha}\cdot\nabla v_{\text{el}}\psi_{R}\rangle=0.$ This gives us the simplified expression in (4.5). Now, using the Cauchy-Schwarz inequality and hypothesis (1.6) one has
$\begin{split}|I|&\leq 2\int|x||\nabla v_{\text{el}}||\psi_{R}||\nabla_{\!\text{A}}\psi_{R}|\\ &\leq 2\varepsilon_{2}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}.\end{split}$
Note that, in the general case, assumption (1.4) is also required to study $I.$ However, this is not necessary here, as $I$ turns out to have the special form (4.5).
Using the bound above and estimating the other terms in (4.4) as in the proof of Theorem 1.2, one has
$\Bigg{(}2-\Big{(}\frac{4d-6}{d-2}\Big{)}\varepsilon_{1}-\Big{(}\frac{4}{d-2}\Big{)}\varepsilon_{2}\,\varepsilon_{3}-\Big{(}\frac{8d-8}{d-2}\Big{)}\varepsilon_{3}-2\varepsilon_{2}-4m\Big{(}\frac{2d-2}{d-2}\Big{)}\varepsilon_{4}\Bigg{)}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}\\ \leq\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},(H_{m}(A,V_{\text{el}})+\lambda)-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle.$
The result follows as in the proof of Theorem 1.2. ∎
Proof of Theorem 1.4.
Since $m=0,$ we can rewrite the eigenvalue equation (3.5) of Lemma 3.1 in a simplified way, namely
$H_{0}(A)^{2}\psi_{R}+2v_{\text{el}}(-i\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}+v_{\text{el}})\psi_{R}-v_{\text{el}}^{2}\psi_{R}-i(\boldsymbol{\alpha}\cdot\nabla v_{\text{el}})\psi_{R}=\lambda^{2}\psi_{R}+(H_{0}(A,V_{\text{el}})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi.$
In the second term of the previous identity we use that $\psi_{R}$ satisfies (3.4), thus we get
$H_{0}(A)^{2}\psi_{R}-i(\boldsymbol{\alpha}\cdot\nabla v_{\text{el}})\psi_{R}=(v_{\text{el}}-\lambda)^{2}\psi_{R}+(H_{0}(A,-V_{\text{el}})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi.$ (4.6)
Multiplying (4.6) by $2(x\cdot\nabla_{\!\text{A}})\psi_{R}+d\psi_{R},$ integrating over $\mathbb{R}^{d}$ and taking the real part of the resulting identity gives
$2\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+2\Im\int x_{k}B_{jk}\psi_{R}^{*}\partial_{j}^{A}\psi_{R}+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla v_{\text{el}})\psi_{R}\rangle=-\int x\cdot\nabla\eta_{\lambda}|\psi_{R}|^{2}\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},(H_{0}(A,-V_{\text{el}})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle,$ (4.7)
where, to simplify the notation, we have defined $\eta_{\lambda}:=(v-\lambda)^{2}.$
Now, multiplying (4.6) by $\psi_{R},$ integrating over $\mathbb{R}^{d}$ and taking the real part of the resulting identity gives
$\begin{split}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+\Re\langle\psi_{R},-&\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi_{R}\rangle+\Re\langle\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla v_{\text{el}})\psi_{R}\rangle\\ &=\int\eta_{\lambda}|\psi_{R}|^{2}+\Re\langle\psi_{R},(H_{0}(A,-V_{\text{el}})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle.\end{split}$ (4.8)
Subtracting (4.8) from (4.7) gives
$\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+2\Im\int x_{k}B_{jk}\psi_{R}^{*}\partial_{j}^{A}\psi_{R}+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+(d-1)\psi_{R},-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+(d-1)\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla v_{\text{el}})\psi_{R}\rangle+\int\eta_{\lambda}|\psi_{R}|^{2}+\int x\cdot\nabla\eta_{\lambda}|\psi_{R}|^{2}\\ =+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+(d-1)\psi_{R},(H_{0}(A,-V_{\text{el}})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle.$ (4.9)
We only need to estimate the terms that depend on $\eta_{\lambda}.$ For the term containing $\nabla\eta_{\lambda}$ we proceed as follows:
$\begin{split}\Big{|}\int x\cdot\nabla\eta_{\lambda}|\psi_{R}|^{2}\Big{|}&\leq\Big{(}\int\frac{|x\cdot\nabla\eta_{\lambda}|^{2}}{\eta_{\lambda}}|\psi_{R}|^{2}\Big{)}^{1/2}\Big{(}\int\eta_{\lambda}|\psi_{R}|^{2}\Big{)}^{1/2}\\ &\leq\Big{(}\int\frac{4|x\cdot\nabla v_{\text{el}}|^{2}(v-\lambda)^{2}}{(v-\lambda)^{2}}|\psi_{R}|^{2}\Big{)}^{1/2}\Big{(}\int\eta_{\lambda}|\psi_{R}|^{2}\Big{)}^{1/2}\\ &\leq 2\Big{(}\int|x|^{2}|\nabla v_{\text{el}}|^{2}|\psi_{R}|^{2}\Big{)}^{1/2}\Big{(}\int\eta_{\lambda}|\psi_{R}|^{2}\Big{)}^{1/2}\\ &\leq\varepsilon_{2}^{2}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+\int\eta_{\lambda}|\psi_{R}|^{2},\end{split}$
where in the last inequality we have used (1.6). Thus, one gets
$\int\eta_{\lambda}|\psi_{R}|^{2}+\int x\cdot\nabla\eta_{\lambda}|\psi_{R}|^{2}\geq-\varepsilon_{2}^{2}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}.$
Estimating the other terms in (4.9) as in the proof of Theorem 1.3 one gets
$\Big{(}1-\Big{(}\frac{4d-7}{d-2}\Big{)}\varepsilon_{1}-2\varepsilon_{2}-\varepsilon_{2}^{2}\Big{)}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}\leq\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+(d-1)\psi_{R},(H_{0}(A,-V_{\text{el}})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle.$
From here the result follows as in the proof of Theorem 1.2. ∎
Proof of Theorem 1.5.
In this specific case, noticing that $\{\boldsymbol{\alpha},V_{\text{sc}}\}=0,$ identity (3.5) in Lemma 3.1 reads as follows
$\Delta_{\text{A}}\psi_{R}-\tfrac{i}{2}\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha}\psi_{R}+(m+v_{\text{sc}})^{2}\psi_{R}-i(\boldsymbol{\alpha}\cdot\nabla v_{\text{sc}})\beta\psi_{R}=\lambda^{2}\psi_{R}+(H_{m}(A,V_{\text{sc}})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R}\psi).$
Multiplying the previous identity by $2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},$ integrating over $\mathbb{R}^{d}$ and taking the real part one gets
$2\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+2\Im\int x_{k}B_{jk}\psi_{R}^{*}\partial_{j}^{A}\psi_{R}+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi_{R}\rangle\\ -\int x\cdot\nabla((v_{\text{sc}}+m)^{2})|\psi_{R}|^{2}+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla v_{\text{sc}})\beta\psi_{R}\rangle\\ =\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},(H_{m}(A,V_{\text{sc}})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle,$ (4.10)
here we have used identity (3.16) in Lemma 3.5. We need to estimate only the first term in the second line of the previous identity as the other terms have been already estimated in the proof of the general result Theorem 1.2. We have
$\begin{split}\Big{|}\int x\cdot\nabla((v_{\text{sc}}+m)^{2})|\psi_{R}|^{2}\Big{|}&\leq 2\int|x||\nabla v_{\text{sc}}||v_{\text{sc}}+m||\psi_{R}|^{2}\\ &\leq 2\int|x||\nabla v_{\text{sc}}||v_{\text{sc}}||\psi_{R}|^{2}+2m\int|x||\nabla v_{\text{sc}}||\psi_{R}|^{2}\\ &\leq 2\Big{(}\int|x|^{2}|\nabla v_{\text{sc}}|^{2}|\psi_{R}|^{2}\Big{)}^{1/2}\Big{(}|v_{\text{sc}}|^{2}|\psi_{R}|^{2}\Big{)}^{1/2}2m\int|x||\nabla v_{\text{sc}}||\psi_{R}|^{2}\\ &\leq 2(\varepsilon_{2}\,\varepsilon_{3}+m\varepsilon_{4}^{2})\|\nabla_{\!\text{A}}\psi_{R}\|^{2},\end{split}$
where in the last inequality we have used (1.6), (1.8) and (1.16). Using the bound above and estimating the other terms in (4.10) as in the proof of Theorem 1.2 one gets
$\Bigg{(}2-\Big{(}\frac{4d-6}{d-2}\Big{)}\varepsilon_{1}-2\varepsilon_{2}\,\varepsilon_{3}-2m\varepsilon_{4}^{2}-\Big{(}\frac{4d-4}{d-2}\Big{)}\varepsilon_{2}\Bigg{)}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}\\ \leq\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},(H_{m}(A,V_{\text{sc}})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle.$
The result follows from the same arguments used in the proof of Theorem 1.2, using condition (1.15). ∎
Before continuing with the proof of Theorem 1.6 we need a preliminary Lemma.
Lemma 4.1.
Let $V_{\text{am}}=i\beta\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}}.$ The following relations hold true
$V_{am}^{2}$ $=|\nabla\phi_{\text{am}}|^{2},$ (4.11)
$\{\beta,V_{\text{am}}\}$ $=0,$ (4.12)
$\{\boldsymbol{\alpha},V_{\text{am}}\}\nabla_{\!\text{A}}$ $=2i\beta(\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}})(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}})+2i\beta\nabla\phi_{\text{am}}\cdot\nabla_{\!\text{A}}$ (4.13)
$(\boldsymbol{\alpha}\cdot\nabla V_{\text{am}})$ $=-i\beta\Delta\phi_{\text{am}}.$ (4.14)
Proof.
The proof follows simply from the commutation relation properties of the Dirac matrices. For the sake of completeness we do the computations. We start by proving (4.11). Note that
$V_{am}^{2}=-(\beta\alpha_{j}\partial_{j}\phi_{\text{am}})(\beta\alpha_{k}\partial_{k}\phi_{\text{am}})=(\alpha_{j}\partial_{j}\phi_{\text{am}})(\alpha_{k}\partial_{k}\phi_{\text{am}})=|\nabla\phi_{\text{am}}|^{2}.$
As for (4.12), one has
$\{\beta,V_{\text{am}}\}=\beta(i\beta\alpha_{j}\partial_{j}\phi_{\text{am}})+i\beta\alpha_{j}\beta\partial_{j}\phi_{\text{am}}=i(\alpha_{j}-\alpha_{j})\partial_{j}\phi_{\text{am}}=0.$
Regarding (4.13), we have
$\begin{split}\{\boldsymbol{\alpha},V_{\text{am}}\}\nabla_{\!\text{A}}&=\{\alpha_{k},i\beta\alpha_{j}\partial_{j}\phi_{\text{am}}\}\partial_{k}^{A}=i\beta(-\alpha_{k}\alpha_{j}+\alpha_{j}\alpha_{k})\partial_{j}\phi_{\text{am}}\partial_{k}^{A}=2i\beta(\alpha_{j}\alpha_{k}+\delta_{j}k)\partial_{j}\phi_{\text{am}}\partial_{k}^{A}\\ &=2i\beta(\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}})(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}})+2i\beta\nabla\phi_{\text{am}}\cdot\nabla_{\!\text{A}}.\end{split}$
As for the last identity (4.14), one has
$(\boldsymbol{\alpha}\cdot\nabla V_{\text{am}})=\alpha_{j}\partial_{j}(i\beta\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}})=-i\beta\alpha_{j}\alpha_{k}\partial_{j}\partial_{k}\phi_{\text{am}}=-i\beta\Delta\phi_{\text{am}}.$
This concludes the proof of the Lemma. ∎
Now we are in position to prove Theorem 1.6.
Proof of Theorem 1.6.
In this specific case, using (4.11)-(4.14) in Lemma 4.1, the identity (3.5) in Lemma 3.1 reads as follows
$H_{m}(A)^{2}\psi_{R}+|\nabla\phi_{\text{am}}|^{2}\psi_{R}-\beta(\Delta\phi_{\text{am}})\psi_{R}+2\beta(\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}})(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}})\psi_{R}+2\beta\nabla\phi_{\text{am}}\cdot\nabla_{\!\text{A}}\psi_{R}\\ =\lambda^{2}\psi_{R}+(H_{m}(A,V_{\text{am}})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R}\psi).$
Multiplying the previous identity by $2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},$ integrating over $\mathbb{R}^{d}$ and taking the real part we get
$2\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+2\Im\int x_{k}B_{jk}\psi_{R}^{*}\partial_{j}^{A}\psi_{R}+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},|\nabla\phi_{\text{am}}|^{2}\psi_{R}\rangle+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-\beta\Delta\phi_{\text{am}}\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},2\beta(\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}})(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi_{R})\rangle+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},2\beta\nabla\phi_{\text{am}}\cdot\nabla_{\!\text{A}}\psi_{R})\rangle\\ =\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},(H_{m}(A,V_{\text{am}})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle.$ (4.15)
We start with the first term in the second line of (4.15). Using (1.18), one has
$\begin{split}|\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},|\nabla\phi_{\text{am}}|^{2}\psi_{R}\rangle|&\leq 2\int|x||\nabla\phi_{\text{am}}|^{2}|\psi_{R}||\nabla_{\!\text{A}}\psi_{R}|+d\int|\nabla\phi_{\text{am}}|^{2}|\psi_{R}|^{2}\\ &\leq 2\varepsilon_{2}^{2}\int\frac{|\psi_{R}|}{|x|}|\nabla_{\!\text{A}}\psi_{R}|+d\varepsilon_{2}^{2}\int\frac{|\psi_{R}|^{2}}{|x|^{2}}\\ &\leq\frac{4}{d-2}\varepsilon_{2}^{2}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+\frac{4d}{(d-2)^{2}}\varepsilon_{2}^{2}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}\\ &=\frac{(8d-8)}{(d-2)^{2}}\varepsilon_{2}^{2}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}.\end{split}$
We continue with the second term in the second line of (4.15), where using (1.19), we have
$\begin{split}|\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-\beta\Delta\phi_{\text{am}}\psi_{R}\rangle|&\leq 2\int|x||\Delta\phi_{\text{am}}||\psi_{R}||\nabla_{\!\text{A}}\psi_{R}|+d\int|\Delta\phi_{\text{am}}||\psi_{R}|^{2}\\ &\leq 2\varepsilon_{3}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+\frac{2d}{d-2}\varepsilon_{3}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}\\ &\leq\frac{4d-4}{d-2}\varepsilon_{3}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}.\end{split}$
As for the first term in the third line of (4.15), one has
$\begin{split}|\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},2\beta(\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}})(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi_{R})\rangle|&\leq 4\int|x||\nabla\phi_{\text{am}}||\nabla_{\!\text{A}}\psi_{R}|^{2}+2d\int|\nabla\phi_{\text{am}}||\psi_{R}||\nabla_{\!\text{A}}\psi_{R}|\\ &\leq 4\varepsilon_{2}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+\frac{4d}{d-2}\varepsilon_{2}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}\\ &\leq\frac{8d-8}{d-2}\varepsilon_{2}\|\nabla_{\!\text{A}}\psi_{R}\|^{2},\end{split}$
where we have used again (1.18). The last term in the third line of (4.15) can be estimated as the previous one.
Plugging the previous estimates in (4.15) we get
$\Bigg{(}2-\Big{(}\frac{4d-6}{d-2}\Big{)}\varepsilon_{1}-\Big{(}\frac{8d-8}{(d-2)^{2}}\Big{)}\varepsilon_{2}^{2}-\Big{(}\frac{4d-4}{d-2}\Big{)}\varepsilon_{3}-2\Big{(}\frac{8d-8}{d-2}\Big{)}\varepsilon_{2}\Bigg{)}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}\\ \leq\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},(H_{m}(A,V_{\text{am}})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle.$
Then the result follows as in the proof of Theorem 1.2, using condition (1.17). ∎
Before proving Theorem 1.7 we need the following Lemma.
Lemma 4.2.
Let $V_{\text{am}}^{3d}=i\beta\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}}-2\beta\boldsymbol{S}\cdot B.$ The following relations hold true
$(V_{am}^{3d})^{2}$ $=|\nabla\phi_{\text{am}}|^{2}+|B|^{2}+2\boldsymbol{\alpha}\cdot(\nabla\phi_{\text{am}}\times B),$ (4.16)
$\{\beta,V_{\text{am}}^{3d}\}$ $=-4\boldsymbol{S}\cdot B,$ (4.17)
$\begin{split}\{\boldsymbol{\alpha},V_{\text{am}}^{3d}\}\nabla_{\!\text{A}}\psi_{R}&=2i\beta(\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}})(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi_{R})+2i\beta\nabla\phi_{\text{am}}\cdot\nabla_{\!\text{A}}\psi_{R}\\ &\phantom{=}-4\beta(\boldsymbol{S}\cdot B)(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}})+2\beta T(B\cdot\nabla_{\!\text{A}}),\end{split}$ (4.18)
$(\boldsymbol{\alpha}\cdot\nabla V_{\text{am}}^{3d})$ $=-i\beta\Delta\phi_{\text{am}}+2\beta\boldsymbol{\alpha}\cdot\nabla(\boldsymbol{S}\cdot B).$ (4.19)
Proof.
We will use the following notation:
$V_{\text{am}}^{3d}=V_{1}+V_{2},\quad\text{where}\quad V_{1}:=i\beta\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}}\quad\text{and}\quad V_{2}:=-2\beta\boldsymbol{S}\cdot B.$
We start proving (4.16). Using (4.11) one has
$V_{am}^{2}=V_{1}^{2}+V_{2}^{2}+\{V_{1},V_{2}\}=|\nabla\phi_{\text{am}}|^{2}+V_{2}^{2}+\{V_{1},V_{2}\}.$
For convenience, we introduce the $4\times 4$ matrix
$T=\begin{pmatrix}0&\mathbb{I}\\ \mathbb{I}&0\end{pmatrix}$.
It is easy to see that for $j=1,2,3$ one has
$2S_{j}T=2TS_{j}=\alpha_{j},\quad\text{or equivalently}\quad\alpha_{j}T=T\alpha_{j}=2S_{j},$ (4.20)
and
$T\beta=-\beta T.$ (4.21)
With (4.20) and (4.21) at hand we can now easily compute $V_{2}^{2}$ and $\{V_{1},V_{2}\}.$ Note that
$V_{2}^{2}=(2\beta S_{j}B_{j})(2\beta S_{k}B_{k})=(\beta T\alpha_{j}B_{j})(\beta T\alpha_{k}B_{k})=(T\alpha_{j}B_{j})(T\alpha_{k}B_{k})=(\boldsymbol{\alpha}\cdot B)(\boldsymbol{\alpha}\cdot B)=|B|^{2}.$
Now,
$\begin{split}\{V_{1},V_{2}\}&=\{i\beta\alpha_{j}\partial_{j}\phi_{\text{am}},-2\beta S_{k}B_{k}\}=iT(\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}})(\boldsymbol{\alpha}\cdot B)-iT(\boldsymbol{\alpha}\cdot B)(\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}})=4iT\boldsymbol{S}\cdot(\nabla\phi_{\text{am}}\times B)\\ &=2\boldsymbol{\alpha}\cdot(\nabla\phi_{\text{am}}\times B),\end{split}$
where we have used the algebraic identity $(\boldsymbol{\alpha}\cdot A)(\boldsymbol{\alpha}\cdot B)=A\cdot B+2i\boldsymbol{S}\cdot(A\times B),$ that holds for any three-component vectors $A,B.$ Gathering these identities gives (4.16). As for (4.17) one has
$\{\beta,V_{\text{am}}^{3d}\}=\{\beta,V_{1}\}+\{\beta,V_{2}\}=0+\{\beta,-2\beta S_{k}B_{k}\}=-4\boldsymbol{S}\cdot B,$
where we have used (4.11) and that for any $k=1,2,3$ one has $\beta S_{k}=S_{k}\beta$, which follows from (4.20) and (4.21). As for (4.18) we have
$\begin{split}\{\boldsymbol{\alpha},V_{\text{am}}^{3d}\}\nabla_{\!\text{A}}&=\{\boldsymbol{\alpha},V_{1}\}\nabla_{\!\text{A}}+\{\boldsymbol{\alpha},V_{2}\}\nabla_{\!\text{A}}=2i\beta(\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}})(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}})+2i\beta\nabla\phi_{\text{am}}\cdot\nabla_{\!\text{A}}+\{\boldsymbol{\alpha},V_{2}\}\nabla_{\!\text{A}}\\ &=2i\beta(\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}})(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}})+2i\beta\nabla\phi_{\text{am}}\cdot\nabla_{\!\text{A}}-4\beta(\boldsymbol{S}\cdot B)(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}})+2\beta T(B\cdot\nabla_{\!\text{A}}).\end{split}$
Here we have used (4.13). Regarding the last identity (4.19),
$(\boldsymbol{\alpha}\cdot\nabla V_{\text{am}}^{3d})=(\boldsymbol{\alpha}\cdot\nabla V_{1})+(\boldsymbol{\alpha}\cdot\nabla V_{2})=-i\beta\Delta\phi_{\text{am}}+(\boldsymbol{\alpha}\cdot\nabla V_{2})=-i\beta\Delta\phi_{\text{am}}+2\beta\boldsymbol{\alpha}\cdot\nabla(\boldsymbol{S}\cdot B).$
where we have used (4.14). This concludes the proof of the Lemma. ∎
Now we are in position to prove Theorem 1.7.
Proof of Theorem 1.7.
In this specific case, using (4.16)-(4.19) in Lemma 4.2, the identity (3.5) in Lemma 3.1 reads as follows
$H_{m}(A)^{2}\psi_{R}+|\nabla\phi_{\text{am}}|^{2}\psi_{R}+|B|^{2}\psi_{R}+2\boldsymbol{\alpha}\cdot(\nabla\phi_{\text{am}}\times B)\psi_{R}-4m\boldsymbol{S}\cdot B\psi_{R}\\ +2\beta(\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}})(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi_{R})+2\beta(\nabla\phi_{\text{am}}\cdot\nabla_{\!\text{A}})\psi_{R}+4i\beta(\boldsymbol{S}\cdot B)(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}})\psi_{R}\\ -2i\beta T(B\cdot\nabla_{\!\text{A}})\psi_{R}-\beta(\Delta\phi_{\text{am}})\psi_{R}-2i\beta(\boldsymbol{\alpha}\cdot\nabla)(\boldsymbol{S}\cdot B)\psi_{R}\\ =\lambda^{2}\psi_{R}+(H_{m}(A,V_{\text{am}}^{3d})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R}\psi).$
Multiplying the previous identity by $2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},$ integrating over $\mathbb{R}^{d}$ and taking the real part we get
$2\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+2\Im\int x_{k}B_{jk}\psi_{R}^{*}\partial_{j}^{A}\psi_{R}+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},|\nabla\phi_{\text{am}}|^{2}\psi_{R}\rangle+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},|B|^{2}\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},2\boldsymbol{\alpha}\cdot(\nabla\phi_{\text{am}}\times B)\psi_{R}\rangle-4m\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},\boldsymbol{S}\cdot B\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},2\beta(\boldsymbol{\alpha}\cdot\nabla\phi_{\text{am}})(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}}\psi_{R})\rangle+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},2\beta(\nabla\phi_{\text{am}}\cdot\nabla_{\!\text{A}})\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},4i\beta(\boldsymbol{S}\cdot B)(\boldsymbol{\alpha}\cdot\nabla_{\!\text{A}})\psi_{R}\rangle+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-2i\beta T(B\cdot\nabla_{\!\text{A}})\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-\beta\Delta\phi_{\text{am}}\psi_{R}\rangle+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-2i\beta(\boldsymbol{\alpha}\cdot\nabla)(\boldsymbol{S}\cdot B)\psi_{R}\rangle\\ =\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},(H_{m}(A,V_{\text{am}}^{3d})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle.$
Proceeding similarly to the proof of Theorem 1.6 and using (1.5), (1.18), (1.19), (1.21) and (1.22), we get
$\Bigg{(}2-\Big{(}\frac{4d-6}{d-2}\Big{)}\varepsilon_{1}-\Big{(}\frac{8d-8}{(d-2)^{2}}\Big{)}\varepsilon_{2}^{2}-\Big{(}\frac{8d-8}{(d-2)^{2}}\Big{)}\varepsilon_{4}^{2}-\Big{(}\frac{16d-16}{(d-2)^{2}}\Big{)}\varepsilon_{2}\,\varepsilon_{4}-m\Big{(}\frac{16d-16}{d-2}\Big{)}\varepsilon_{1}\\ -\Big{(}\frac{16d-16}{d-2}\Big{)}\varepsilon_{2}-\Big{(}\frac{24d-24}{d-2}\Big{)}\varepsilon_{4}-\Big{(}\frac{4d-4}{d-2}\Big{)}\varepsilon_{3}-\Big{(}\frac{8d-8}{d-2}\Big{)}\varepsilon_{5}\Bigg{)}\|\nabla_{\!\text{A}}\psi_{R}\|^{2}\\ \leq\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},(H_{m}(A,V_{\text{am}}^{3d})+\lambda)(-i\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle.$
Acknowledgements
N. A. was supported by the grants PID2021-126813NB-I00, funded by MCIN/AEI/10.13039/501100011033, and IT1615-22, funded by the Basque Government.
L.C. was supported by the grant Ramón y Cajal RYC2021-032803-I funded by MCIN/AEI/10.13039/50110 and by Ikerbasque.
M.M was supported by the grant PID2021-126813NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by the FSE+.
Data availability
There is no data associated with this article
Conflict of interest
There is no conflict of interest to declare